OpenStudy (anonymous):

example of an addition problem that simplifies to 4x

OpenStudy (anonymous):

4x + 0

OpenStudy (anonymous):

2x+2x 3x+1x

OpenStudy (campbell_st):

you can have 2x + 2x = 4x x + 3x = 4x 5x + (-x) = 4x

OpenStudy (anonymous):

328+4x-328 :)